# MonSter: Marry Monodepth to Stereo Unleashes Power

### v1

@@ 4.3 Performance in Ill-posed Regions
MonSter fully leverages the advantages of monocular depth priors, effectively overcoming challenges in ill-posed regions. To validate this, we conduct comparisons on several representative ill-posed regions, such as reflective areas, edge and non-edge regions, and distant backgrounds:
@@
MonSter(Ours) 5.66 6.81 2.75 3.38 1.73 2.13
Table 3: Results of the reflective regions on KITTI 2012 leaderboard. The best and second best are marked with colors.
@@
MonSter(Ours) 1.91 18.59 0.31 3.57
Table 4: Comparison on SceneFlow test set in different regions.
  • Reflective Regions: We conducted comparisons on the reflective regions of KITTI 2012 benchmark. MonSter ranks $1^{st}$ on the KITTI 2012 leaderboard for all metrics of reflective regions. As shown in Tab.3, we have elevated the existing SOTA to a new level. MonSter surpasses Selective-IGEV and LoS by 30.16% and 29.24% on Out-4(All) metric, respectively. Notably, compared to the SOTA method NMRF-Stereo, we achieved significant improvements of 58.32% and 65.02% in the Out-3(All) and Out-4(All) metrics respectively.
@@
  • Distant Backgrounds: Stereo matching struggles with depth perception for distant objects, we improve it by incorporating our SGA and MGR modules. As shown in Tab.2, MonSter improves the D1-bg metric by 18.12% compared to the baseline on KITTI 2015 benchmark. The D1-bg metric reflects the percentage of stereo disparity outliers averaged specifically over background regions.
Method KITTI-12 KITTI-15 Middlebury ETH3D
($>$3px) ($>$3px) ($>$2px) ($>$1px)
@@
CFNet[29] 4.72 5.81 9.80 5.80
RAFT-Stereo[20] 5.12 5.74 9.36 3.28
CREStereo[16] 5.03 5.79 12.88 8.98
Selective-IGEV[35] 5.64 6.05 12.04 5.40
NMRF-Stereo[13] 4.23 5.10 7.54 3.82
IGEV[38] 4.84 5.51 6.23 3.62
MonSter(Ours) 3.62 3.97 5.17 2.03
Training Set Scene Flow + CREStereo + Tartan Air
IGEV[38] 3.95 4.30 5.56 2.38
MonSter(Ours) 2.95 3.23 2.94 1.21
Table 5: Zero-shot generalization benchmark. The top half is generalization comparisons only trained on Scene Flow. The bottom half compares generalizations for models trained on a combination of Scene Flow, CREStereo, and TartanAir datasets.
4.4 Zero-shot Generalization
Our method not only improves the accuracy of in-domain datasets but also enhances generalization ability. We evaluate the generalization performance of MonSter from synthetic datasets to unseen real-world scenes. All models were trained on the same synthetic datasets and then tested directly on real-world datasets: KITTI[11, 23], Middlebury[27], and ETH3D[28] training sets. We conducted a more thorough generalization comparison by training with different datasets:
Scene Flow: We compared with SOTA methods which are only trained on the Scene Flow training set. As shown in the top half of Tab. 5, MonSter achieves the best generalization performance across four datasets. MonSter even surpasses CFNet and CREStereo, both of which are specifically designed for cross-domain generalization.
@@ 4.4 Zero-shot Generalization
Figure 6: Visual comparisons on our captured real-world zero-shot data provide a more comprehensive assessment of generalization capability. All models are trained solely on the synthetic SceneFlow dataset. MonSter significantly outperforms IGEV in textureless regions, reflective areas and fine structures, etc.
Mix 3 datasets: Interestingly, only adding CREStereo and TartanAir to the training set, the generalization of MonSter improves significantly. This is because the SceneFlow dataset differs substantially from real-world distributions, whereas the virtual data in TartanAir more closely resemble real-world scenes. This similarity helps our SGA and MGR modules better in learning how to effectively integrate monocular and stereo cues. As shown in the bottom half of Tab. 5, we surpass the baseline[38] by 47.12% and 49.16% on Middlebury and ETH3D respectively.
The strong sim-to-real generalization of MonSter is highly encouraging. In the future, we plan to scale up the size and diversity of simulation training data to further boost the performance of MonSter.
@@
Full model (MonSter) MGR SGA 0.37 4.25 0.64
Table 6: Ablation study of the effectiveness of proposed modules on the Scene Flow test set.
@@
Table 7: Efficiency and universality of MonSter.
4.5 Ablation Study
To demonstrate the effectiveness of each component of MonSter, we ablation on Scene Flow [22] and the following strategies are discussed:
  • Disparity Fusion: Monocular depth often carries substantial noise, which complicates its fusion with the stereo branch’s disparity. To demonstrate the efficiency of our Mono Guided Refinement (MGR) fusion method, we replaced the MGR module with a convolution-based hourglass network of equal parameter numbers, directly concatenating the monocular and stereo disparities for fusion, denoted as Mono+Conv. As shown in Tab.6, compared with ‘Mono+Conv’, ‘Mono+MGR’ improves the EPE metric by 6.52%.
[…]
  • Feature Sharing: The feature extraction component of monocular depth estimation models, pre-trained on large datasets, contains rich contextual information that can be shared with the stereo branch. We employ a feature transfer network to map these features, transforming them into a representation compatible with stereo matching. As shown in the last 2 rows of Tab.6, feature sharing improved the EPE metric by 5.13%.
Compatibility with Other Monocular Models. To validate the generalizability of MonSter to different depth models, we replace the monocular branch with the ViT-large version of [45] and the dpt_beit_large version of MiDaS [25]. As shown in Tab.7, all MonSter variants consistently outperform the baseline [38]. This demonstrates the versatility of MonSter and its potential to benefit from future advancements in monocular depth models.
@@ 4.5 Ablation Study
Efficiency. We use the ViT-large version of DepthAnythingV2 [46] as the monocular branch, which introduces time and memory overheads. Compared to the baseline [38], our inference time increases from 0.37s to 0.64s. With the monocular branch containing 335.3M parameters, the stereo branch having 12.6M parameters, and the SGA and MGR modules having 8.2M parameters, our full model has a total of 356.1M parameters. Existing methods [1, 36] also use ViT-based encoders with a large number of parameters. For example, CroCo-Stereo [36] has 437.4M parameters, larger than MonSter, yet its performance is inferior, as shown in Tab.2, this demonstrates that the number of parameters does not determine the final effectiveness. Our research focuses on the accuracy and generalization capabilities, given the significant improvement in accuracy and generalization (as shown in Tab.2 and Tab.5), the additional inference cost is acceptable. Notably, our approach simplifies stereo matching and enables us to achieve SOTA performance with only 4 iterations ($N_{1}$=$N_{2}$=2). As shown in Tab.7, while the baseline requires 32 iterations, our method achieves a 10.64% improvement in EPE with just 4 iterations, achieving a better accuracy-speed trade-off. Further memory reduction can be achieved through encoder quantization or distillation, which we leave as future work.
5 Conclusion
We propose MonSter, a novel pipeline that decouples the stereo matching task into a simpler paradigm of recovering scale and shift from relative depth. This approach fully leverages the contextual and geometric priors provided by monocular methods while avoiding issues such as noise and scale ambiguity. As a result, our method significantly enhances the accuracy and robustness of stereo matching in ill-posed regions. MonSter achieves substantial improvements over the state-of-the-art on five widely used benchmarks. Additionally, MonSter also demonstrates the best generalization performance, and experimental results show that there is still considerable room for improvement. Therefore, our future work will focus on scaling up MonSter to serve as a stereo foundation model, enabling a wide range of downstream applications that require metric depth.

### v2

@@ IV-D Performance in Ill-posed Regions
MonSter++ fully leverages the advantages of monocular depth priors, effectively overcoming challenges in ill-posed regions. To validate this, we conduct comparisons on several representative ill-posed regions, such as reflective areas, edge and non-edge regions, and distant backgrounds:
TABLE IV: Quantitative comparisons on the KITTI 2012 [39] and KITTI 2015 [23] stereo matching benchmarks. Our RT-MonSter++ runs 4 update iterations at inference. The best and second best are marked with colors.
Method KITTI 2012 [39] KITTI 2015 [23]
Run-time
Out-2 Out-2 Out-3 Out-3 Out-4 Out-4 Out-5 Out-5 D1-fg D1-all D1-fg D1-all
Noc All Noc All Noc All Noc All Noc Noc All All
[…]
DeepPrunerFast [106] - - - - - - - - 3.43 2.35 3.91 2.59 50
AANet [28] 2.30 2.96 1.55 2.04 1.20 1.58 0.98 1.30 3.66 1.85 3.96 2.03 60
BGNet+ [107] 2.78 3.35 1.62 2.03 1.16 1.48 0.90 1.16 3.76 2.01 4.09 2.19 35
CoEx [108] 2.54 3.09 1.55 1.93 1.15 1.42 0.91 1.13 3.09 1.86 3.41 2.02 27
Fast-ACVNet+ [109] 2.39 2.97 1.45 1.85 1.06 1.36 0.85 1.09 3.29 1.85 3.53 2.01 45
[…]
TABLE V: Results of the reflective regions on KITTI 2012 leaderboard. The best and second best are marked with colors.
@@
[…]
TABLE VI: Comparison on SceneFlow test set in different regions. The best result is bolded.
@@
MonSter++ (Ours) 1.91 18.59 0.31 3.57
  • Reflective Regions: We conducted comparisons on the reflective regions of KITTI 2012 benchmark. MonSter++ ranks $1^{st}$ on the KITTI 2012 leaderboard for all metrics of reflective regions. As shown in Tab.V, we have elevated the existing SOTA to a new level. MonSter++ surpasses Selective-IGEV and LoS by 24.92% and 23.92% on Out-4 (All) metric, respectively. Notably, compared to the SOTA method NMRF-Stereo, we achieved significant improvements of 57.46% and 62.40% on the Out-3 (All) and Out-4 (All) metrics respectively.
@@
  • Distant Backgrounds: Stereo matching struggles with depth perception for distant objects, and we improve it by incorporating our SGA and MGR modules. As shown in Tab.II, MonSter++ improves the D1-bg metric by 18.84% compared to the baseline on KITTI 2015 benchmark. The D1-bg metric reflects the percentage of stereo disparity outliers averaged specifically over background regions.
TABLE VII: Zero-shot generalization benchmark for real-time stereo matching methods and High-accuracy stereo matching methods. The best and second best are marked with colors.
Target Method KITTI-12 KITTI-15 ETH3D Middlebury Avg. Runtime
(>>3px) (>>3px) (>>1px) (>>2px) (ms)
Real-Time
@@
DeepPrunerFast [106] 16.8 15.9 11.0 30.8 18.63 50
BGNet+ [107] 24.8 20.1 10.3 37.0 23.05 35
CoEx [108] 13.5 11.6 9.0 25.5 14.90 27
Fast-ACVNet+ [109] 12.4 10.6 7.9 20.1 12.75 45
[…]
Accuracy
Training Set Scene Flow
[…]
CREStereo[30] 5.0 5.8 9.0 12.9 8.18 410
Selective-IGEV[5] 5.6 6.1 5.4 12.0 7.28 240
[…]
IGEV++[33] 6.2 6.4 4.1 7.8 6.13 280
[…]
Mix-Training
Training Set Full Training Set (FTS)
RT-IGEV++[33] 3.9 4.6 2.5 8.3 4.83 48
[…]
IV-E Real-Time Version of MonSter++
To enable practical deployment on mobile platforms such as UAVs and robots, we design a real-time lightweight version of MonSter++ for the stereo matching task, named RT-MonSter++. Benefiting from MonSter++’s effective exploitation of monocular priors and a carefully designed coarse-to-fine strategy, RT-MonSter++ achieves over 20 FPS inference speed at 1K resolution while significantly surpassing the accuracy of state-of-the-art methods. Following IGEV++ [33], we conduct inference at the KITTI resolution of $1248\times 384$ and compare our RT-MonSter++ with state-of-the-art real-time approaches on KITTI 2012 and KITTI 2015 benchmarks. As shown in Tab. IV, RT-MonSter++ achieves the best performance among all published methods. On the KITTI 2012 benchmark, RT-MonSter++ surpasses the previous best method, RT-IGEV++ [33], by 19.23% and 20.75% on the Out-4 (ALL) and Out-5 (ALL) metrics, respectively. On the KITTI 2015 benchmark, RT-MonSter++ outperforms CoEx [108] and RT-IGEV++ [33] by 19.42% and 21.45% on D1-fg (Noc) metric, respectively.
Benefiting from the advantages of monocular priors, RT-MonSter++ also achieves the best generalization performance among real-time methods. We evaluate its generalization ability on the half-resolution training set of the Middlebury 2014 dataset [21], the ETH3D [20] training set, and the training sets of KITTI 2012 [39] and KITTI 2015 [23]. As shown in Tab. VII, our RT-MonSter++ achieves state-of-the-art performance in the zero-shot setting among real-time stereo matching methods. Especially on the challenging Middlebury dataset with large disparities, RT-MonSter++ improves upon the previous best method, RT-IGEV++, by reducing the 2-pixel outlier rate from 17.3 to 11.4, corresponding to a 34.1% improvement.
It is worth highlighting that RT-MonSter++ demonstrates the strongest scalability among real-time methods. When trained on the Full Training Set (FTS), RT-MonSter++ achieves remarkable zero-shot performance gains over RT-IGEV++ across all four benchmarks, with particularly significant improvements of 32.0% on ETH3D and 25.3% on Middlebury. Moreover, RT-MonSter++ delivers performance comparable to the accuracy-oriented IGEV model, while requiring only one-third of its inference cost. We will release RT-MonSter++ model that trained on over 2 million image pairs to the community to empower this highly efficient and generalizable method for real-world deployment of depth perception on mobile platforms.
TABLE VIII: Zero-shot generalization performance on DrivingStereo [110] under different weather. We utilize >>3px All in comparisons. The best and second best are marked with colors.
Method Sunny Cloudy Rainy Foggy Avg.
[…]
@@ IV-E Real-Time Version of MonSter++
Figure 8: Visual comparisons on our captured real-world zero-shot data provide a more comprehensive assessment of generalization capability. All models are trained solely on the synthetic SceneFlow dataset. MonSter++ significantly outperforms IGEV in textureless regions, reflective areas and fine structures, etc.
IV-F Zero-shot Generalization
Our method not only improves the accuracy of in-domain datasets but also enhances the generalization ability. We evaluate the generalization performance of MonSter++ from synthetic datasets to unseen real-world scenes. All models were trained on the same synthetic datasets and then tested directly on real-world datasets: KITTI[39, 23], Middlebury[21], and ETH3D[20] training sets. We conducted a more thorough generalization comparison by training with different datasets: 1) Scene Flow: We compared with SOTA methods which are only trained on the Scene Flow training set. As shown in Tab. VII, MonSter++ achieves the best generalization performance across four datasets in accuracy models. Compared to our baseline method IGEV, we reduce the average error across four datasets from 5.03 to 3.70, achieving a 26.4% improvement. MonSter++ even surpasses CFNet and CREStereo, both of which are specifically designed for cross-domain generalization. Fig. 4 highlights MonSter++’s improvements in reflective, textureless, and fine-structured regions. 2) Full Training Set: To demonstrate the scalability of our method, we further compare its generalization performance with the baseline method when trained on the Full Training Set. As shown in the Mix Training setting in Tab. VII, our MonSter++ surpass the baseline method IGEV[6] by 39.5% and 60.0% on Middlebury and ETH3D, respectively. Notably, under the same mix-training setting, RT-MonSter++ delivers performance comparable to the accuracy-oriented IGEV model while requiring only 47 ms per inference—approximately one-third of IGEV’s runtime (180 ms).
DrivingStereo: In addition, we compare our method with state-of-the-art approaches specifically designed for generalization on the DrivingStereo[110] dataset. We choose this dataset for evaluation because existing generalization methods often rely on heterogeneous training data, such as incorporating private datasets or synthesizing right-view images, which makes direct comparison less fair. In contrast, the DrivingStereo dataset is unseen during training for all methods and contains diverse real-world scenarios with challenging weather conditions, providing a fair benchmark to assess a model’s ability to generalize to the real world. As shown in Tab. VIII, MonSter++ achieves the best overall performance across all weather conditions. In particular, on rainy scenes, it reduces the 3px error rate from 6.08 to 3.08 compared with the previous best method SMoEStereo [112], yielding a remarkable 49.3% improvement. Furthermore, compared with the state-of-the-art FoundationStereo, MonSter++ lowers the average error across four weather conditions from 4.93 to 2.69, demonstrating its outstanding generalization ability in challenging environments.
TABLE IX: Ablation study of the effectiveness of proposed modules on the Scene Flow test set. All runtime results are measured under the same setting of 32 iterations. The best result is bolded.
@@ IV-F Zero-shot Generalization
Full model (MonSter++) MGR SGA 0.37 4.25 0.64
IV-G Ablation Study
To demonstrate the effectiveness of each component of MonSter++, we ablation on Scene Flow [22] and the following strategies are discussed:
  • Disparity Fusion: Monocular depth often carries substantial noise, which complicates its fusion with the stereo branch’s disparity. To demonstrate the efficiency of our Mono Guided Refinement (MGR) fusion method, we replaced the MGR module with a convolution-based hourglass network of equal parameter numbers, directly concatenating the monocular and stereo disparities for fusion, denoted as Mono+Conv. As shown in Tab.IX, compared with ‘Mono+Conv’, ‘Mono+MGR’ improves the EPE metric by 6.52%.
[…]
  • Feature Sharing: The feature extraction component of monocular depth estimation models, pre-trained on large datasets, contains rich contextual information that can be shared with the stereo branch. We employ a feature transfer network to map these features, transforming them into a representation compatible with stereo matching. As shown in the last 2 rows of Tab.IX, feature sharing improved the EPE metric by 5.13%.
Compatibility with Other Monocular Models. To validate the generalizability of MonSter++ to different depth models, we replace the monocular branch with the ViT-large version of [66] and the dpt_beit_large version of MiDaS [113]. As shown in Tab.X, all MonSter++ variants consistently outperform the baseline [6]. This demonstrates the versatility of MonSter++ and its potential to benefit from future advancements in monocular depth models.
TABLE X: Efficiency and universality of MonSter++. The best result is bolded.
@@ IV-G Ablation Study
Efficiency. We use the ViT-large version of DepthAnythingV2 [38] as the monocular branch, which introduces time and memory overheads. Compared to the baseline [6], our inference time increases from 0.37s to 0.64s. With the monocular branch containing 335.3M parameters, the stereo branch having 12.6M parameters, and the SGA and MGR modules having 8.2M parameters, our full model has a total of 356.1M parameters. Existing methods [114, 90] also use ViT-based encoders with a large number of parameters. For example, CroCo-Stereo [90] has 437.4M parameters, larger than MonSter++, yet its performance is inferior, as shown in Tab.II, this demonstrates that the number of parameters does not determine the final effectiveness. Our research focuses on the accuracy and generalization capabilities, given the significant improvement in accuracy and generalization (as shown in Tab.II, Tab.VII and Tab.VIII), the additional inference cost is acceptable. Notably, our approach simplifies stereo matching and enables us to achieve SOTA performance with only 4 iterations ($N_{1}$$N_{2}$=2). As shown in Tab.X, while the baseline requires 32 iterations, our method achieves a 10.64% improvement in EPE with just 4 iterations, achieving a better accuracy-speed trade-off. Further memory reduction can be achieved through encoder quantization or distillation, which we leave as future work.
V Conclusion
We propose MonSter++, a novel pipeline that decouples the multi-view depth estimation into a simpler paradigm of recovering scale and shift from relative depth. This approach fully leverages the contextual and geometric priors provided by monocular methods while avoiding issues such as noise and scale ambiguity. Consequently, our approach markedly improves both the accuracy and robustness of depth estimation, particularly in ill-posed regions. MonSter++ achieves substantial gains over state-of-the-art methods across eight widely used benchmarks spanning three tasks—stereo matching, real-time stereo matching, and multi-view stereo. We also introduce a real-time version, RT-MonSter++, which delivers the best accuracy among all published real-time methods. To foster highly generalizable depth models for the open-source community, we conduct large-scale training on 2 million image pairs for both MonSter++ and RT-MonSter++. All models and training code will be publicly released.
Acknowledgement. Authors from HUST are supported by National Natural Science Foundation of China(62472184) and the Fundamental Research Funds for the Central Universities.